OpenStudy (anonymous):

How many solutions does this equation have? 3(4x - 3) - 7x = 5x -9 A. One solution B. infinite solutions C. No solutions

OpenStudy (texaschic101):

3(4x - 3) - 7x = 5x - 9 --- distribute through the parenthesis 12x - 9 - 7x = 5x - 9 -- combine terms 5x - 9 = 5x - 9 0 = 0 when you get an answer such as this...0 = 0 or 5 = 5 or -2 = -2......this means there can be any number subbed in for x and it would make the equation true

OpenStudy (texaschic101):

and if you can use any number for x and it would make the equation true, what does that tell you about the equation

OpenStudy (anonymous):

She's right, I'm sorry, It's been a year since I did one of these problems. A bit rusty.

OpenStudy (texaschic101):

it happens...if you don't do it enough. I, unfortunately, do not have a life, so I am on here doing lots of these...so sad

OpenStudy (anonymous):

Yeah, I've been doing honors geometry this year, so I lost some things.

OpenStudy (texaschic101):

ohhhh...I hate geometry....algebra is ok

OpenStudy (anonymous):

I like geometry.

OpenStudy (texaschic101):

you should do real good on OS then....tons of geometry q's

OpenStudy (anonymous):

OS?

OpenStudy (texaschic101):

openstudy...

OpenStudy (anonymous):

Oh duh. Honestly I havent been here since right before summer vacation.

OpenStudy (anonymous):

Take this to the message system instead?

OpenStudy (texaschic101):

well.....that calls for a WELCOME BACK :)

OpenStudy (anonymous):

Thanks!

OpenStudy (texaschic101):

I got some stuff to do so I gotta go now....maybe I will see you around...well..not really see, but you know what I mean :)

OpenStudy (anonymous):

Cool, send me a message later when you get back on, make Hermes proud!

OpenStudy (texaschic101):

lol...sure thing :) catch ya later

OpenStudy (anonymous):

Bye!